# PITT-PACC-2408

### v1

@@ Abstract
The interpretation of groups of particle spins at colliders as quantum states has opened up the possibility of using colliders for quantum information. While most efforts have focused on utilizing the decays of the particles to infer their spins to reconstruct the quantum density matrix, we show that the production kinematics of the particles provides sufficient information about the spins to establish quantum tomography without using the decays. We perform a comparative study, highlighting the advantages and disadvantages of using this “kinematic approach” relative to the usual “decay approach.” Since the kinematic approach leverages the simplicity of scattering kinematics, this approach promises to achieve the optimal statistical results for quantum tomography at colliders.
@@ I Introduction
Figure 1: Schematic illustration for a quantum state described by the density matrix $\rho$ at the center, that may be probed by the production kinematics or by the angular variables of the decay products.
The ways that we understand the microscopic world are based on quantum mechanics. Two notable phenomena that quantum systems can exhibit – which classical systems cannot – are entanglement and Bell inequality violation. Recently these phenomena have been studied at high energy colliders where the spins of the particles produced compromise the quantum state [1, 2, 3]. For instance, in $t\bar{t}$ production at the Large Hadron Collider (LHC), the $t$ and $\bar{t}$ are treated as qubits leading to a bipartite qubit system [4, 5, 6, 7, 8, 9, 10]. Depicting the full properties of this quantum state is known as quantum tomography.
Traditionally, quantum tomography is a procedure that involves performing a complementary set of quantum measurements on an ensemble of identically-prepared quantum states. Different measurements collapse the states differently and provide complementary information about the quantum state. At colliders, typically spin is used as the qubit. The spins of particles produced at high-energy colliders are not directly measured but are inferred from the directions of decay products. A second distinctive feature of quantum information at colliders is that “fictitious states” [5, 11, 12] – states which are dependent on the choice of spin quantization axis [11, 12] – are reconstructed rather than genuine quantum states.
@@ II Two Approaches
With this perspective in mind, we show that there are two natural sets of parameters to describe the density matrix. We label the use of the first set of parameters as the decay approach. This is the standard technique in past works [1, 2, 3]. In this Letter, we introduce the other set of parameters and label their usage as the kinematic approach, which is a distinct method. This only depends on the $2\to 2$ production kinematics and offers unique advantages to quantum tomography at colliders. These approaches are drawn schematically in Fig. 1. The essential quantum information, contained in $\rho$, may be probed either via the kinematics of production or via the angular distributions of decay products.
@@ II.1 The Decay Approach
The statistical uncertainty $\Delta C_{ij}$ on the measurement of an entry of the spin correlation matrix $C_{ij}$ is universal
$\Delta C_{ij}=\frac{3}{\sqrt{N}},$ (8)
where $N$ is the number of observed events.
@@
$\Delta\mathcal{C}$ $=\frac{1}{2}\sqrt{(\Delta C_{11})^{2}+(\Delta C_{22})^{2}+(\Delta C_{33})^{2}},$ (9)
$\Delta\mathcal{B}$ $=\sqrt{2}\sqrt{(\Delta C_{22})^{2}+(\Delta C_{33})^{2}}.$ (10)
Anticipating improvements in future measurements, we can define the percentage accuracy for these measurements as
$\frac{\Delta\mathcal{C}}{\mathcal{C}}$ $=\frac{3\sqrt{3}}{(C_{11}+C_{33}-C_{22}-1)}\frac{1}{\sqrt{N}},$ (11)
$\frac{\Delta\mathcal{B}}{\mathcal{B}}$ $=\frac{3\sqrt{2}}{C_{33}-C_{22}}\frac{1}{\sqrt{N}}.$ (12)
@@ II.2 The Kinematic Approach
where $c_{\Theta}=\cos\Theta$ and $s_{\Theta}=\sin\Theta$. The concurrence $\mathcal{C}$ can thus be measured event-by-event and follows a distribution $F_{\mathcal{C}}(\Theta,\beta)$. The distribution is obtained as
$F_{\mathcal{C}}(\Theta,\beta)$ $=\frac{C_{11}(\Theta,\beta)+C_{33}(\Theta,\beta)-C_{22}(\Theta,\beta)-1}{2},$ (15)
@@
The measured Bell variable follows a distribution
$F_{\mathcal{B}}(\Theta,\beta)$ $=\sqrt{2}|C_{33}(\Theta,\beta)-C_{22}(\Theta,\beta)|$ (16)
@@
The measured values are
$\mathcal{C}=\langle F_{\mathcal{C}}(\Theta,\beta)\rangle,\quad\mathcal{B}=\langle F_{\mathcal{B}}(\Theta,\beta)\rangle.$ (17)
The statistical uncertainties of concurrence $\Delta\mathcal{C}$ and the Bell variable $\Delta\mathcal{B}$ depend on the distributions
$\Delta\mathcal{C}=\frac{\sqrt{\text{Var}(F_{\mathcal{C}})}}{\sqrt{N}},\quad\Delta\mathcal{B}=\frac{\sqrt{\text{Var}(F_{\mathcal{B}})}}{\sqrt{N}},$ (18)
where ${\rm Var}(F_{\mathcal{R}})=\langle(\mathcal{R}-\mathcal{R}(\Theta,\beta))^{2}\rangle$ where $\mathcal{R}(\Theta,\beta)$ is the “random variable” and $\mathcal{R}$ is the mean. The explicit forms of $F_{\mathcal{C}}$ and $F_{\mathcal{B}}$ are process-dependent.
@@
$\frac{\Delta\mathcal{C}}{\mathcal{C}}\leq\frac{1}{2\langle F_{\mathcal{C}}(\Theta,\beta)\rangle}\frac{1}{\sqrt{N}},\ \ \frac{\Delta\mathcal{B}}{\mathcal{B}}\leq\frac{1}{\langle F_{\mathcal{B}}(\Theta,\beta)\rangle}\frac{1}{\sqrt{N}}.\hfill$ (19)
Compared with Eqs. (11) and (12), the statistical uncertainties in the kinematic approach are substantially smaller than in the decay approach. This is because the kinematic approach exploits the full kinematic dependence in measuring quantities while the decay approach infers the spin correlation statistically. For a fair assessment of the experimental precision, one needs to account for systematic uncertainties, especially when the number of events is sufficiently large.
@@ III.1 Quantum Tomography in Drell-Yan Production $pp\to Z\to\mu^{+}\mu^{-}$
Figure 2: Complete quantum tomography distribution for $Z\to\mu^{+}\mu^{-}$ at the LHC. All parameters of Eq. (1) that are not shown are zero.
The spin correlation matrix of $\mu^{+}\mu^{-}$ produced from a $Z$ boson decay is given by Eq. (32). We obtain the concurrence and Bell variable distributions as
$F_{\mathcal{C}}(\Theta,\beta\approx 1)$ $=\frac{s_{\Theta}^{2}(g_{A}^{2}-g_{V}^{2})}{(1+c_{\Theta}^{2})(g_{A}^{2}+g_{V}^{2})},$ (20)
$F_{\mathcal{B}}(\Theta,\beta\approx 1)$ $=\frac{2\sqrt{2}\left(g_{A}^{2}+g_{V}^{2}c_{\Theta}^{2}\right)}{\left(1+c_{\Theta}^{2}\right)\left(g_{A}^{2}+g_{V}^{2}\right)},$ (21)
where $g_{V}$ and $g_{A}$ are, respectively, the vector and axial coupling strengths of the $Z$ boson to the charged leptons. [^3] The couplings of $Z$ to initial state quarks and the collision energies are irrelevant for $pp$ collisions; see Appendix A for details. $\theta_{W}$ is the Weinberg angle.
We show the full quantum tomography for $Z\to\mu^{+}\mu^{-}$ at the LHC in Fig. 2 and calculate the concurrence and Bell variable in Fig. 3. In Fig. 3(a) for the angular distributions of $\mathcal{C}$ and $\mathcal{B}$, we see that the events with the largest quantum effects are in the central region. This is because the $\mu^{+}\mu^{-}$ system is mostly a spin-triplet with larger entanglement for $J_{z}=0$. In Fig. 3(b), we calculate the integrated state with a selection cut of $|\cos\Theta|<c_{\rm cut}$ to better isolate the signals of entanglement and Bell non-locality. The horizontal dashed lines denote the boundaries of Bell inequality violation and entanglement.
Figure 3: (a) Quantum tomography distribution for $Z\to\mu^{+}\mu^{-}$ at the LHC, (b) concurrence $\mathcal{C}$ and the Bell variable $\mathcal{B}$ integrated over angles satisfying $|\cos\Theta|<c_{\rm cut}$. The horizontal dashed lines indicate the boundaries of the entanglement and Bell non-locality.
The experimental observability is determined by the reachable uncertainties $\Delta\mathcal{C}$ and $\Delta\mathcal{B}$, which depends on the statistics as given in Eq. (18). The variance, over the full phase space, is
${\rm Var}(F_{\mathcal{B}})=2{\rm Var}(F_{\mathcal{C}})=\frac{3(\pi-3)(g_{A}^{2}-g_{V}^{2})^{2}}{4(g_{A}^{2}+g_{V}^{2})^{2}}.$ (22)
The LHC has a large number of $Z\to\mu^{+}\mu^{+}$ events [13] which will accumulate $10^{8}$ events from a luminosity of $300~{}{\rm fb}^{-1}$ at Run 3. The percentage accuracies, including only statistical uncertainty, of $\Delta\mathcal{C}/\mathcal{C}$ and $\Delta\mathcal{B}/\mathcal{B}$ can easily achieve $0.01\%$. At this level of precision, even a sub-percent level systematic uncertainty would dominate the error budget.
@@ III.2 Quantum Tomography in $pp\to t\bar{t}$
where $L_{q\bar{q}}$ and $L_{gg}$ are the parton luminosities, respectively, of $q\bar{q}$ and $gg$. Unlike in the process $pp\to Z\to\mu^{+}\mu^{-}$, the distributions $F_{\mathcal{C}}(\Theta,\beta)$ and $F_{\mathcal{B}}(\Theta,\beta)$ depend on the parton luminosities and require a numerical calculation.
@@
Where the difference between approaches arises is in the statistical uncertainties on the $\mathcal{C}$ and $\mathcal{B}$ measurements. These uncertainties are calculated in Table 1 at the LHC with an integrated luminosity of 300 fb-1. For both approaches, we include the fully leptonic $t\bar{t}$ branching ratio and the only selection cuts we apply are $m_{t\bar{t}}<350~{}{\rm GeV}$ in the threshold region and $m_{t\bar{t}}>1.5~{}{\rm TeV}$ and $|\cos\Theta|<0.5$. The resulting statistical uncertainties using the decay approach are $1\%-26\%$. In all cases the kinematic approach achieves an uncertainty that is more than an order of magnitude smaller. [^4] We do not include reconstruction efficiencies. While the absolute uncertainties may change in a fully realistic simulation, the ratio between approaches is robust.
In practice, these two approaches rely on different observables and will have different systematic uncertainties. The decay approach requires parameter estimation through fits which generally increases the uncertainty relative to theory-level predictions. The kinematic approach, on the other hand, is largely determined by momentum measurements. We thus claim that the kinematic approach yields the optimal statistical sensitivity.
[…]
@@
Table 1: Statistical uncertainties on $\mathcal{C}$ and $\mathcal{B}$ measurements for $pp\to t\bar{t}$ with representative selection cuts. An integrated luminosity of 300 fb-1 is used. The fully leptonic branching ratio of $t\bar{t}$ is included in both approaches without additional kinematic cuts.
@@ IV Discussion and Conclusions
Although we illustrated our results with the two-qubit system, the kinematic approach is also applicable to other quantum states, like qubit-qutrit systems like $tW$, or two-qutrit systems like $W^{+}W^{-}$. This is because the kinematic observables for unpolarized $2\to 2$ production are fully determined by the speed and direction of motion of the outgoing particles in the center-of-mass frame. This approach can be generalized to an $n$-body quantum system as well, in terms of the minimal number of kinematic variables.
The kinematic approach involves the fewest kinematic observables to construct the predicted density matrix, while the decay approach measures more angular variables to infer the spin information. The two approaches are thus complementary and offer meaningful cross-checks of the same quantum information. This can be useful, for example, if there exist beyond-the-Standard-Model predictions that modify the production and decay of the qubit particles differently.
Due to the simpler measurement and less demanding event reconstruction, the kinematic approach offers tremendous upside in the reconstruction of the quantum density matrix at colliders. With the density matrix, many quantum quantities, for example quantum discord [16], EPR steering [16], or magic [17], can be calculated. Furthermore, this method is sensitive to the effects from new physics beyond the Standard Model that may couple to the quantum state at production, such as a new pseudoscalar – a CP-odd Higgs $A$ or a $t\bar{t}$ bound state $\eta_{t}$ – or a vector – a $Z^{\prime}$ vector boson or a $t\bar{t}$ bound state $\Upsilon_{t}$ – to modify the spin-singlet or spin-triplet predictions.
In conclusion, we have proposed a new approach to explore the quantum information at colliders, the kinematic approach, which is based on $2\to 2$ production kinematics. There is significant complementarity between this approach and the decay approach which uses particle decays to compute spin correlations. Due to the simplicity of $2\to 2$ kinematics and particle reconstruction, the kinematic approach promises to achieve the optimal statistical outcome for quantum tomography of the relevant processes and has a wide range of applicability. In particular, we can move beyond the scope of using only promptly-decaying particles to using nearly the full array of observable particles in the Standard Model to construct the quantum information.

### v2

@@ Abstract
In constructing quantum states at colliders, most efforts have focused on utilizing the particle decays to infer the spins to reconstruct the quantum density matrix. We show that the production kinematics of the particles provides sufficient information about the spins to establish quantum tomography event-by-event without using the decays. We lay out the advantages of using this “kinematic approach” relative to the usual “decay approach.” Since the kinematic approach leverages the simplicity of scattering kinematics, this approach promises to achieve optimal statistical results for quantum tomography at colliders.
@@ I Introduction
The ways that we understand the microscopic world are based on quantum mechanics. Two notable phenomena that quantum systems can exhibit – which classical systems cannot – are entanglement and Bell inequality violation. Recently these phenomena have been studied at high energy colliders where the quantum states are the spin states of the particle system [1, 2, 3]. For instance, in $t\bar{t}$ production at the Large Hadron Collider (LHC), the $t$ and $\bar{t}$ are treated as qubits leading to a bipartite qubit system [4, 5, 6, 7, 8, 9, 10]. Depicting the full properties of this quantum state is known as quantum tomography.
Traditionally, quantum tomography is a procedure that involves performing a complementary set of quantum measurements on an ensemble of identically-prepared quantum states. Different measurements collapse the states differently and provide complementary information about the quantum state. At colliders, typically the spin state of a particle is used as the qubit. The spins of particles produced at high-energy colliders are not directly measured but are inferred from the directions of decay products. A second distinctive feature of quantum information at colliders is that “fictitious states” [5, 11, 12] are reconstructed when using event-dependent basis, rather than genuine quantum states when using a fixed basis.
@@ II Two Approaches
With this perspective in mind, we show that there are two natural sets of parameters to describe the density matrix. We label the use of the first set of parameters as the decay approach. This is the standard technique in past works [1, 2, 3]. In this Letter, we introduce the other set of parameters and label their usage as the kinematic approach, which is a distinct method. This only depends on the $2\to 2$ production kinematics and offers unique advantages to quantum tomography at colliders. The essential quantum information, contained in $\rho$, may be probed either via the kinematics of production or via the angular distributions of decay products.
@@ II.1 The Decay Approach
From Eq. (4), the statistical uncertainty $\Delta C_{ij}^{{\rm stat}}$ on the measurement of an entry of the spin correlation matrix $C_{ij}$, which follows the normal distribution for a large number of events $N$, takes the universal form of
$\Delta C_{ij}^{{\rm stat}}=\frac{3}{\sqrt{N}}.$ (8)
@@
$\Delta\overline{\mathcal{C}}^{\rm stat}$ $=\frac{1}{2}\sqrt{(\Delta C_{11}^{{\rm stat}})^{2}+(\Delta C_{22}^{{\rm stat}})^{2}+(\Delta C_{33}^{{\rm stat}})^{2}},$ (9)
$\Delta\overline{\mathcal{B}}^{{\rm stat}}$ $=\sqrt{2}\sqrt{(\Delta C_{22}^{{\rm stat}})^{2}+(\Delta C_{33}^{{\rm stat}})^{2}}.$ (10)
where the statistical uncertainties of different $C_{ij}$’s are approximately independent [13]. Anticipating improvements in future measurements, we can define the percentage accuracy for these measurements as
$\frac{\Delta\overline{\mathcal{C}}^{{\rm stat}}}{\overline{\mathcal{C}}}$ $=\frac{3\sqrt{3}}{(C_{11}+C_{33}-C_{22}-1)}\frac{1}{\sqrt{N}},$ (11)
$\frac{\Delta\overline{\mathcal{B}}^{{\rm stat}}}{\overline{\mathcal{B}}}$ $=\frac{3\sqrt{2}}{C_{33}-C_{22}}\frac{1}{\sqrt{N}}.$ (12)
@@ II.2 The Kinematic Approach
where $c_{\Theta}=\cos\Theta$ and $s_{\Theta}=\sin\Theta$. The concurrence $\mathcal{C}$ can thus be measured event-by-event from the two kinematic variables $\Theta$ and $\beta$:
$\mathcal{C}(\Theta,\beta)$ $=\frac{C_{11}(\Theta,\beta)+C_{33}(\Theta,\beta)-C_{22}(\Theta,\beta)-1}{2},$ (15)
@@
The Bell variable is
$\mathcal{B}(\Theta,\beta)$ $=\sqrt{2}|C_{33}(\Theta,\beta)-C_{22}(\Theta,\beta)|$ (16)
@@
The measured values of concurrence and the Bell variable are, respectively, the mean of the variables $\mathcal{C}$ and $\mathcal{B}$
$\overline{\mathcal{C}}=\langle\mathcal{C}(\Theta,\beta)\rangle,\quad\overline{\mathcal{B}}=\langle\mathcal{B}(\Theta,\beta)\rangle.$ (17)
With sufficient large number of events, the mean values of $\mathcal{C}(\Theta,\beta)$ and $\mathcal{B}(\Theta,\beta)$ follow normal distributions and the statistical uncertainties of concurrence $\Delta\mathcal{C}$ and the Bell variable $\Delta\mathcal{B}$ are
$\Delta\overline{\mathcal{C}}^{{\rm stat}}=\frac{\sqrt{\text{Var}(\mathcal{C}(\Theta,\beta))}}{\sqrt{N}},\quad\Delta\overline{\mathcal{B}}^{{\rm stat}}=\frac{\sqrt{\text{Var}(\mathcal{B}(\Theta,\beta))}}{\sqrt{N}},$ (18)
where ${\rm Var}(\mathcal{R})=\langle(\overline{\mathcal{R}}-\mathcal{R})^{2}\rangle$, $\mathcal{R}$ is the “random variable” and $\overline{\mathcal{R}}$ is its mean value. The explicit forms of $\mathcal{C}$ and $\mathcal{B}$ regarding kinematic variables are process-dependent.
@@
$\frac{\Delta\overline{\mathcal{C}}^{{\rm stat}}}{\overline{\mathcal{C}}}\leq\frac{1}{2\overline{\mathcal{C}}}\frac{1}{\sqrt{N}},\ \ \frac{\Delta\overline{\mathcal{B}}^{{\rm stat}}}{\overline{\mathcal{B}}}\leq\frac{\sqrt{2}}{\overline{\mathcal{B}}}\frac{1}{\sqrt{N}}.\hfill$ (19)
Compared with Eqs. (11) and (12), the statistical uncertainties in the kinematic approach are substantially smaller than in the decay approach. This is because the kinematic approach exploits the full kinematic dependence in measuring quantities while the decay approach infers the spin correlation statistically.
@@ III.1 Quantum Tomography in Drell-Yan Production $pp\to Z\to\mu^{+}\mu^{-}$
The spin correlation matrix of $\mu^{+}\mu^{-}$ produced from a $Z$ boson decay is given by Eq. (32). We obtain the concurrence and the Bell variable as
$\mathcal{C}(\Theta,\beta\approx 1)$ $=\frac{s_{\Theta}^{2}(g_{A}^{2}-g_{V}^{2})}{(1+c_{\Theta}^{2})(g_{A}^{2}+g_{V}^{2})},$ (20)
$\mathcal{B}(\Theta,\beta\approx 1)$ $=\frac{2\sqrt{2}\left(g_{A}^{2}+g_{V}^{2}c_{\Theta}^{2}\right)}{\left(1+c_{\Theta}^{2}\right)\left(g_{A}^{2}+g_{V}^{2}\right)},$ (21)
where $g_{V}$ and $g_{A}$ are, respectively, the vector and axial coupling strengths of the $Z$ boson to the charged leptons, [^3] See Appendix A for the calculation of the variables $\mathcal{B}$, $\mathcal{C}$ and their variances. and $\Theta$ is the scattering angle between the outgoing $\mu^{-}$ and the incoming beam in the $Z$ boson rest frame.
We show the full quantum tomography for $Z\to\mu^{+}\mu^{-}$ at the LHC and calculate the concurrence and Bell variable in Fig. 1. In (a) for the angular distributions of $\mathcal{C}$ and $\mathcal{B}$, we see that the events with the largest quantum effects are in the central region. This is because the $\mu^{+}\mu^{-}$ system is mostly a spin-triplet with larger entanglement for $J_{z}=0$. In (b), we calculate the integrated state with a selection cut of $|\cos\Theta|<c_{\rm cut}$ to better isolate the signals of entanglement and Bell non-locality. The horizontal dashed lines denote the boundaries of Bell inequality violation and entanglement.
Figure 1: (a) Quantum tomography distribution for $Z\to\mu^{+}\mu^{-}$ at the LHC, (b) concurrence $\overline{\mathcal{C}}$ and the Bell variable $\overline{\mathcal{B}}$ integrated over angles satisfying $|\cos\Theta|<c_{\rm cut}$. The horizontal dashed lines indicate the boundaries of the entanglement and Bell non-locality.
The experimental observability is determined by the reachable uncertainties $\Delta\overline{\mathcal{C}}^{\rm stat}$ and $\Delta\overline{\mathcal{B}}^{{\rm stat}}$, which depends on the statistics as given in Eq. (18). The variance, over the full phase space, is
${\rm Var}(\mathcal{B})=2{\rm Var}(\mathcal{C})=\frac{3(\pi-3)(g_{A}^{2}-g_{V}^{2})^{2}}{2(g_{A}^{2}+g_{V}^{2})^{2}}.$ (22)
The LHC has a large number of $Z\to\mu^{+}\mu^{+}$ events [14] which will accumulate $10^{8}$ events from a luminosity of $300~{}{\rm fb}^{-1}$ at Run 3. The percentage accuracies, including only statistical uncertainty, of $\Delta\overline{\mathcal{C}}^{\rm stat}/\overline{\mathcal{C}}$ and $\Delta\overline{\mathcal{B}}^{\rm stat}/\overline{\mathcal{B}}$ can easily achieve $0.01\%$.
@@ III.2 Quantum Tomography in $pp\to t\bar{t}$
where $L_{q\bar{q}}$ and $L_{gg}$ are the parton luminosities, respectively, of $q\bar{q}$ and $gg$. Unlike in the process $pp\to Z\to\mu^{+}\mu^{-}$, the variables $\mathcal{C}(\Theta,\beta)$ and $\mathcal{B}(\Theta,\beta)$ depend on the parton luminosities and require a numerical calculation.
@@
Where the difference between approaches arises is in the statistical uncertainties on the $\overline{\mathcal{C}}$ and $\overline{\mathcal{B}}$ measurements. The statistical uncertainties are calculated in Table 1 at the LHC with an integrated luminosity of 300 fb-1. For both approaches, we include the fully leptonic $t\bar{t}$ branching ratio and the only selection cuts we apply are $m_{t\bar{t}}<350~{}{\rm GeV}$ in the threshold region and $m_{t\bar{t}}>1.5~{}{\rm TeV}$ and $|\cos\Theta|<0.5$. The resulting statistical uncertainties using the decay approach are $1\%-26\%$. In all cases the kinematic approach achieves a statistical uncertainty that is more than an order of magnitude smaller. [^4] We do not include reconstruction efficiencies. While the absolute uncertainties may change in a fully realistic simulation, the ratio between approaches is robust.
cuts $\overline{\mathcal{C}}$ $\Delta\overline{\mathcal{C}}^{\rm stat}$$\overline{\mathcal{B}}$ $\Delta\overline{\mathcal{B}}^{\rm stat}$
@@
Table 1: Statistical uncertainties on $\overline{\mathcal{C}}$ and $\overline{\mathcal{B}}$ measurements for $pp\to t\bar{t}$ with representative selection cuts. An integrated luminosity of 300 fb-1 is used. The fully leptonic branching ratio of $t\bar{t}$ is included in both approaches without additional kinematic cuts.
@@ IV Discussion and Conclusions
Although we illustrated our results with the two-qubit system, the kinematic approach is also applicable to other quantum states such as two qutrit system $W^{\pm}$, as the spin density matrix is also determined by the two kinematic variable $\Theta$ and $\beta$ [17]. This approach can be generalized to an $n$-body quantum system as well, in terms of the minimal number of kinematic variables.
The kinematic approach also involves the fewer kinematic observables for particle pairs such as $t\bar{t}$ that decay. This is because the kinematic approach only needs the total momentum of top decay products while the decay approach requires additional angular separations between top decay products. The two approaches are thus complementary and offer meaningful cross-checks of the same quantum information. This can be useful, for example, if there exist beyond-the-Standard-Model predictions that modify the production and decay of the particles differently, then the decay approach and kinematic approach are sensitive to different modifications.
Due to the simpler measurement and less demanding event reconstruction, the kinematic approach offers tremendous upside in reconstructing the quantum density matrix at colliders. With the density matrix, many quantum quantities, for example quantum discord [18, 19], EPR steering [18], or magic [20], can be calculated. Furthermore, this method is sensitive to the effects from new physics beyond the Standard Model that may couple to the quantum state at production, such as a new pseudoscalar – a CP-odd Higgs $A$ or a $t\bar{t}$ bound state $\eta_{t}$ – or a vector – a $Z^{\prime}$ vector boson or a $t\bar{t}$ bound state $\Upsilon_{t}$ – to modify the spin-singlet or spin-triplet predictions.
Our presentation has focused on statistical uncertainties. In realistic projections, systematic uncertainty should be included. Unless an experiment has sufficient data that their results are systematics dominated, reducing the statistical uncertainty is of critical importance. In a study of low energy $e^{+}e^{-}\to\tau^{+}\tau^{-}$ scattering, it was found that the reduced statistical uncertainty plays a crucial role in the observability even considering realistic systematic uncertainties [21].
Throughout our presentation, we illustrate our method for quantities that are linear with respect to the spin correlation matrix elements $C_{ij}$. For non-linear cases, one should first measure the mean and variance of the spin correlation matrix elements in accordance with the fictitious states, and then construct the non-linear variables, similar to the decay method.
In conclusion, we have proposed a new approach to explore the quantum information at colliders, the kinematic approach, which is based on $2\to 2$ production kinematics. There is significant complementarity between this approach and the decay approach which uses particle decays to compute spin correlations. Due to the simplicity of $2\to 2$ kinematics and particle reconstruction, the kinematic approach has a smaller statistical uncertainty. It also benefits from easily including multiple decay channels leading to an effectively larger data sample, and thus achieves the optimal statistical outcome for quantum tomography of the relevant processes and has a wide range of applicability. In particular, we can move beyond the scope of using only promptly-decaying particles to using nearly the full array of observable particles in the Standard Model to construct the quantum information.